# A Customized SAT-based Solver for Graph Coloring

### v1

@@ Abstract
We leverage the recently introduced IPASIR-UP interface for CaDiCal to implement these techniques with a SAT solver. Furthermore, we propose new features that take advantage of the underlying SAT solver. These include modifying the integrated decision strategy with vertex domination hints and using incremental bottom-up search that allows to reuse learned clauses from previous calls. Additionally, we integrate a more efficient clique computation to improve the lower bounds during the search.
We validate the effectiveness of each new feature through an experimental analysis. ZykovColor outperforms other state-of-the-art graph coloring implementations on the DIMACS benchmark set. Further experiments on random Erdős-Rényi graphs show that our new approach dominates state-of-the-art SAT-based methods for both very sparse and highly dense graphs.
@@ 1 Introduction
Graph coloring consists of assigning a minimum number of colors to all vertices of a graph $G$ such that no two adjacent vertices are assigned the same color. The graph coloring problem is to find the chromatic number $\chi(G)$, i.e., the smallest number of colors needed for such an assignment. Graph coloring is a classical combinatorial optimization problem with many applications [8, 15, 24] that is notoriously difficult to solve. It is NP-hard to approximate $\chi(G)$ within a factor $n^{1-\epsilon}$ for all $\epsilon>0$ [37].
@@
Recently, the IPASIR-UP interface [12] for the SAT solver CaDiCal [4] was introduced. It allows to modify the course of the SAT solver by registering callbacks to all major steps of the algorithm, including custom propagation, adding clauses, decisions and lazy reasoning.
@@ 1.1 Our Contribution
In this work, we adapt the propagator for the graph coloring problem presented in [19]. Originally proposed in conjunction with a hybrid CP/SAT approach, we adopt it to be used with the state-of-the-art SAT solver CaDiCal [4], heavily using the recently introduced IPASIR-UP interface [12] (Sections 3.1, 3.2 and 3.3). We use the features of the interface to present ZykovColor, a customized SAT-based solver for the graph coloring problem.
The new algorithmic ideas introduced in this paper include
  • Implementing custom constraint propagations utilizing the IPASIR-UP interface with lazy (deferred) addition of reason clauses (Section 3.1),
  • using better clique heuristics for lower bound pruning (Section 3.4),
  • a new decision strategy, where we use vertex domination hints to modify the decisions made by the SAT solver (Section 3.5),
  • performing an incremental bottom-up search for the chromatic number without re-initializing the solver for each decision problem (Section 3.6).
We perform an ablation study demonstrating the effectiveness of each feature in ZykovColor (Section 4.3). Additionally, we perform extensive computational studies, benchmarking ZykovColor with several state-of-the-art algorithms (Section 4.4). On the DIMACS benchmarks for graph coloring, ZykovColor outperforms other state-of-the-art algorithms. Additionally, we analyze the performance of different algorithms across varying graph densities using a large set of Erdős-Rényi graphs. Finally, with a little more computation time, ZykovColor successfully colors the previously unsolved DIMACS instance wap07 (Section 4.5).
The paper starts by introducing the graph coloring problem, the traditional assignment encoding and the full Zykov encoding, as well as the IPASIR-UP interface [12] in Section 2.
@@ 2 Preliminaries
Let $G=(V,E)$ be an undirected graph with vertices $V$ and edges $E$. In usual notation, $n\coloneqq|V|$, $m\coloneqq|E|$, and $N_{G}(u)=\{v\in V|\{u,v\}\in E\}$ is the set of neighbors of $u$, shortened to $N(u)$ if the graph is clear from the context.
We assume that the vertices are numbered from 1 to $n$, i.e. $V=\{1,\dots,n\}$. This allows us to give short precise descriptions of certain encodings later. A coloring of size $k$ or $k$-coloring is a map $f:V\to\{1,\ldots,k\}$ such that $f(u)\neq f(v)$ for all edges $\{u,v\}\in E$. The graph coloring problem is to compute a coloring using the minimum possible number of colors; the size of such a coloring is the chromatic number denoted as $\chi(G)$.
@@
We solve the graph coloring problem using a Satisfiability (SAT) solver, which takes a formula on Boolean variables in conjunctive normal form as input. It returns as the solution either SAT with a satisfying variable assignment or UNSAT if no satisfying assignment exists. We will consider formulas that decide the $k$-coloring problem, i.e., that are satisfiable if and only if a $k$-coloring exists. Consequently, we must solve a series of decision problems to determine the smallest $k$ such that the graph is $k$-colorable but not $(k-1)$-colorable.
This outer iteration might motivate to use a MaxSAT solver instead. However, we only consider SAT solvers for several reasons: After determining lower and upper bounds in preprocessing, usually only a few decision problems need to be solved. We also observed that for most instances, only the two problems $k\in\{\chi(G)-1,\chi(G)\}$ take a significant time. Finally, we want to interact with the solving process. For the SAT solver CaDiCal [4] we can use the IPASIR-UP interface [12]. We are not aware of similar interfaces in competitive MaxSAT solvers.
@@ 2.1 Assignment Encoding
A natural satisfiability encoding to decide $k$-colorability is the assignment encoding. It is called traditional encoding in [34]. It uses $n\cdot k$ variables $x_{vi}\in\{\operatorname{True},\operatorname{False}\}$, one for each vertex $v\in V$ and color $i\in\{1,\ldots,k\}$. The variable $x_{vi}$ is true exactly if vertex $v$ is assigned color $i$. The following clauses assert that a satisfiable assignment corresponds to a $k$-coloring, while proving UNSAT shows that the graph has no $k$-coloring.
@@
Equations 1 and 2 ensure that each vertex is assigned exactly one color, while Equation 3 enforces that adjacent vertices get different colors.
@@ 2.2 Zykov-based Encoding
An alternative encoding is based on the Zykov recurrence 4 for the chromatic number [38].
$\chi(G)=\min\{\chi(G/(u,v),\chi(G+(u,v)))\}\quad\text{ if }\{u,v\}\notin E$ (4)
In an optimal coloring, two non-adjacent vertices will either have the same color or different ones. These two cases are simulated by either merging the vertices or by adding an edge. An example of a Zykov recursion step of a graph is given in Figure 1. Recursive application yields a so-called Zykov tree, where all leaf vertices represent complete graphs without non-adjacent vertices. Any leaf with $k$ vertices determines a $k$-coloring for the original graph by assigning the same color to all vertices that were merged on the path from the root to the leaf in the Zykov tree. Therefore, any leaf with a minimum number of vertices determines an optimal coloring and the chromatic number of the original graph. The Zykov recursion was used successfully in branch-and-price frameworks [29, 26, 20].
[…]
Figure 1: Root node and first two child nodes of Zykov tree for the graph $C_{5}$.
In conjunction with a SAT solver it was first used in [33] to break symmetries on top of the assignment encoding 1, 3 and 2. Later, the satisfiability encoding that we will discuss next became the basis of [16] and [19].
For each non-edge $\{u,v\}\notin E$, we define Boolean variables $e_{uv}$, where $e_{uv}=\operatorname{True}$ means that $u$ and $v$ are merged, and thus have the same color. Likewise, $e_{uv}=\operatorname{False}$ means that the edge $\{u,v\}$ is added to the graph. For ease of notation, we extend the variables $e_{uv}$ to proper edges $\{u,v\}\in E$. They are forced to $e_{uv}=\operatorname{False}$.
To ensure that each (partial) assignment of truth values corresponds to a valid graph in the Zykov tree, transitivity constraints (5) are required. They assert that if $u,v$ and $v,w$ have the same color, vertices $u$ and $w$ must also have the same color. With just the clauses from (5), any satisfying assignment corresponds to a leaf node in the Zykov tree but the number of colors is not yet restricted. To this end, Glorian et al. [16] use auxiliary variables $c_{v}$ for each vertex $v\in V$. In the clauses (6), $c_{v}=\operatorname{True}$ means that vertex $v$ needs a new color different from any previous vertex in the order given by $V=\{1,\dots,n\}$. Now, the sum over the $c_{v}$ variables counts the total number of used colors. The at-most-$k$-constraints (7) can be modeled as SAT in multiple ways, e.g.  using cardinality networks [1] or the totalizer encoding [2]. This yields the following full (Zykov) encoding for the $k$-coloring problem.
@@ 2.3 IPASIR-UP Interface
First, we briefly repeat the main steps in a CDCL solver:
@@
The external propagator is notified by the SAT solver of newly assigned literals, new decision levels, and backtracking. It then has the possibility to modify the solving process using callbacks, as visualized in Figure 2. For our application, we will use the following functionalities of the interface:
@@ 3 SAT-based propagator for the Zykov encoding
In this section we adopt concepts of the propagator presented by Hébrard and Katsirelos  [19] to the satisfiability setting using the interface discussed in Section 2.3. This includes details on the transitivity constraints, using lower bounds for search tree pruning, and additional inferred propagations (Sections 3.1, 3.2 and 3.3).
We further extend the SAT implementation with novel ideas and improvements, namely a better clique computation, modifications to the decision strategy, and the design of an incremental bottom-up optimization that can reuse clauses and learned information (Sections 3.4, 3.5 and 3.6).
@@ 3.1 Transitivity constraints
We follow the approach in [19] to avoid adding the cubic number of transitivity clauses (5) of the full encoding. Instead, the propagator ensures that $e_{uw}$ is set to true whenever $e_{uv}$ and $e_{vw}$ are set to true. This is done efficiently by maintaining a graph $H$ that corresponds to the graph $G^{\prime}$, where $G^{\prime}$ is the graph in the Zykov tree obtained by following the merge operations and edge additions given by the currently assigned variables $e_{uv}$ at that point of the search.
@@
These propagations are handled via cb_propagate, and the explanation clauses (5) are provided lazily using cb_add_reason, only should they be needed during conflict analysis. This is an advantage over the CP/SAT hybrid in [19], where transitivity clauses are immediately added to the encoding when the propagation is given. An important observation is that the propagator considers each variable only once per search path. This way, the transitivity propagation from the root node down to a full assignment only has complexity $O(n^{2})$[19]. In contrast, the constraint propagation from the root to a full assignment based on the full (cubic) encoding of all transitivity constraints (5) has cubic complexity.
@@ 3.2 Lower bound for pruning
This lower bound-based pruning also enforces that at most $k$ colors are used [19]. If we find a lower bound that is larger than $k$, the assignment is pruned and the search is backtracked. This way, we do not need to add the clauses 6 and 7 explicitly. This also avoids the auxiliary $c_{v}$ variables in the encoding.
@@
To prune a partial assignment, we need to add a clause to the encoding that is unsatisfied under the current assignment. Since both the clique and Mycielskian bounds are based on a subgraph $H$, the explanation can be given by a description of the subgraph that needs to be avoided:
@@
The cb_add_external callback is used to add the clause to the encoding. Because it is unsatisfied at the node where the subgraph is present and the lower bound is found, this causes the desired backtracking. The clique heuristic is called at every node of the search tree, while the Mycielskian lower bound is only computed after backtracking and if, the gap between clique and upper bound is at most one. We did not observe a significant gain in pruning when the Mycielskian lower bound is computed regardless of the gap, in particular for bottom-up optimization. We thus choose to call the heuristic less frequently.
@@ 3.3 Inferred propagations
In this section, we discuss another strategy. It does not prune the current partial assignment but infers necessary variable assignments. Given a subgraph with chromatic number $k$ and vertices $u$ and $v$, the propagator can determine that $e_{uv}$ must be assigned either true or false, if otherwise a subgraph with chromatic number $k+1$ is formed.
@@
Given a subgraph $H=(V^{\prime},E^{\prime})$ in $G=(V,E)$ with chromatic number $k$, if there exists a vertex $v\in V\backslash V^{\prime}$ such that $V^{\prime}\backslash N(v)=\{u\}$ then $G+(u,v)$ has chromatic number at least $k+1$.
@@
Given a subgraph $H=(V^{\prime},E^{\prime})$ in $G=(V,E)$ with chromatic number $k$, if there exist two vertices $u$ and $v$ such that $V^{\prime}\subseteq(N(u)\cup N(v))$ then $G/(uv)$ has chromatic number at least $k+1$.
@@ 3.4 Improved clique computation
While very fast, the clique heuristic used in [19] (see Section Section 3.2) is a greedy algorithm. It often does not produce near-optimum cliques. To improve the pruning potential of the propagator, we prefer an algorithm that produces larger cliques. Since the clique algorithm is called frequently in the propagator, keeping its running time small is essential.
@@ 3.5 Improved decision strategy
We use the reduction rule of dominated vertices to determine ?good? decisions and communicate them to the SAT solver. A vertex $u$ is dominated by a vertex $v$ if $N(u)\subseteq N(v)$ and we know that vertex $u$ can always take the same color as the more constrained vertex $v$. We can therefore merge $u$ into $v$ as there exists an optimal solution where the two vertices have the same color. Note that we can potentially also find an optimal solution where $u$ and $v$ have different colors. Thus, domination is weaker than the inferred propagations from the previous section. Therefore, we provide $e_{uv}=\operatorname{True}$ as a decision for the SAT solver using the cd_decide callback. The idea is that following this decision will lead to a part in the search tree that is more likely to contain a solution and thus should be checked first.

### v2

@@ Abstract
We leverage the recently introduced IPASIR-UP interface for CaDiCaL to implement these techniques with a SAT solver. Furthermore, we propose new features that take advantage of the underlying SAT solver. These include modifying the integrated decision strategy with vertex domination hints and using incremental bottom-up search that allows to reuse learned clauses from previous calls. Additionally, we integrate a more effective clique computation and an algorithm for computing the fractional chromatic number to improve the lower bounds used for pruning during the search.
We validate the effectiveness of each new feature through an experimental analysis. ZykovColor outperforms other state-of-the-art graph coloring implementations on the DIMACS benchmark set. Further experiments on random Erdős-Rényi graphs show that our new approach matches or outperforms state-of-the-art SAT-based methods for both very sparse and highly dense graphs. We give an additional configuration of ZykovColor that dominates other SAT-based methods on the Erdős-Rényi graphs.
@@ 1 Introduction
Graph coloring consists of assigning a minimum number of colors to all vertices of a graph $G$ such that no two adjacent vertices are assigned the same color. The graph coloring problem is to find the chromatic number $\chi(G)$, i.e., the smallest number of colors needed for such an assignment. It is a classical combinatorial optimization problem with many applications [7, 15, 24] that is notoriously difficult to solve. It is NP-hard to approximate $\chi(G)$ within a factor $n^{1-\epsilon}$ for all $\epsilon>0$ [38].
@@
SAT solvers were employed successfully not only for exact coloring. In recent work, a generic local search SAT solver found colorings matching known lower bounds for the hard DIMACS instance wap07a [8].
Recently, the IPASIR-UP interface [12] for the SAT solver CaDiCaL [4] was introduced. It allows to modify the course of the SAT solver by registering callbacks to all major steps of the algorithm, including custom propagation, adding clauses, decisions, and lazy reasoning.
@@ 1.1 Our Contribution
In this work, we adapt the propagator for the graph coloring problem presented in [19]. Originally proposed in conjunction with a hybrid CP/SAT approach, we adopt it to be used with the state-of-the-art SAT solver CaDiCaL [4], heavily using the recently introduced IPASIR-UP interface [12] (Sections˜3.1, 3.2 and 3.3). We use the features of the interface to present ZykovColor, a customized SAT-based solver for the graph coloring problem.
Our new algorithms and their engineering include:
  • Implementing custom constraint propagations utilizing the IPASIR-UP interface with lazy (deferred) addition of reason clauses (Section˜3.1),
  • using better clique heuristics for lower bound pruning (Section˜3.4),
  • a new decision strategy, where we use vertex domination hints to modify the decisions made by the SAT solver (Section˜3.5),
  • performing an incremental bottom-up search for the chromatic number without re-initializing the solver for each decision problem (Section˜3.6),
  • using the fractional chromatic number to compute stronger lower bounds for more effective pruning on dense graphs (Section˜3.7).
We perform an ablation study demonstrating the effectiveness of each feature in ZykovColor (Section˜4.3). Additionally, we perform extensive computational studies, benchmarking ZykovColor with several state-of-the-art algorithms (Section˜4.4). On the DIMACS benchmarks for graph coloring, ZykovColor outperforms other state-of-the-art algorithms. Additionally, we analyze the performance of different algorithms across varying graph densities using a large set of Erdős-Rényi graphs.
@@ 2 Preliminaries
Let $G=(V,E)$ be an undirected graph with vertices $V$ and edges $E$. In usual notation, $n\coloneqq|V|$, $m\coloneqq|E|$, and $N_{G}(u)=\{v\in V|\{u,v\}\in E\}$ is the set of neighbors of $u$, shortened to $N(u)$ if the graph is clear from the context. The (edge) density of $G$ is the value $\frac{2m}{n(n-1)}$.
We assume that the vertices are numbered from 1 to $n$, i.e., $V=\{1,\dots,n\}$. This allows us to give short precise descriptions of certain encodings later. A coloring of size $k$ or $k$-coloring is a map $f:V\to\{1,\ldots,k\}$ such that $f(u)\neq f(v)$ for all edges $\{u,v\}\in E$. The graph coloring problem is to compute a coloring using the minimum possible number of colors; the size of such a coloring is the chromatic number, denoted $\chi(G)$.
@@
To solve the graph coloring problem, we employ a satisfiability (SAT) solver that takes a Boolean formula in conjunctive normal form as input. It returns as the solution either SAT with a satisfying variable assignment or UNSAT if no satisfying assignment exists. We will consider formulas that decide the $k$-coloring problem, i.e., that are satisfiable if and only if a $k$-coloring exists. Consequently, we must solve a series of decision problems to determine the smallest $k$ such that the graph is $k$-colorable but not (k1)(k-1)-colorable.
This outer iteration might motivate the use of a MaxSAT solver instead. However, we only consider SAT solvers for several reasons: After determining lower and upper bounds in preprocessing, usually only a few decision problems need to be solved. We also observed that for most instances, only the two problems $k\in\{\chi(G)-1,\chi(G)\}$ take a significant time. Finally, we want to interact with the solving process. For the SAT solver CaDiCaL [4] we can use the IPASIR-UP interface [12]. We are not aware of similar interfaces in competitive MaxSAT solvers.
@@ 2.1 Assignment Encoding
A natural satisfiability encoding to decide $k$-colorability is the assignment encoding. It is called traditional encoding in [35]. It uses $n\cdot k$ variables $x_{vi}\in\{\operatorname{True},\operatorname{False}\}$, one for each vertex $v\in V$ and color $i\in\{1,\ldots,k\}$. The variable $x_{vi}$ is $\operatorname{True}$ exactly if vertex $v$ is assigned color $i$. The following clauses assert that a satisfiable assignment corresponds to a $k$-coloring, while proving UNSAT shows that the graph has no $k$-coloring.
@@
Equations˜1 and 2 ensure that each vertex is assigned exactly one color, while Equation˜3 enforces that adjacent vertices get different colors.
@@ 2.2 Zykov-based Encoding
An alternative encoding is based on the following Zykov recurrence for the chromatic number [39].
$\chi(G)=\min\{\chi(G/(u,v)),\chi(G+(u,v))\}\;\text{ if }\{u,v\}\notin E$ (4)
In an optimal coloring, two non-adjacent vertices $u,v$ will either have the same color or different ones. These two cases are simulated by either merging the vertices ($G/(u,v)$) or by adding an edge between them ($G+(u,v)$). An example of a Zykov recursion step for a graph is given in Figure˜1. Recursive application yields a so-called Zykov tree, where all leaf vertices represent complete graphs. Any leaf with $k$ vertices determines a $k$-coloring for the original graph by assigning the same color to all vertices that were merged on the path from the root to the leaf in the Zykov tree. Therefore, any leaf with a minimum number of vertices determines an optimal coloring and the chromatic number of the original graph. The Zykov recursion was used successfully in branch-and-price frameworks [30, 27, 20].
Figure 1: Root node and first two child nodes of the Zykov tree for the graph $C_{5}$.
In conjunction with a SAT solver it was first used in [34] to break symmetries on top of the assignment encoding Equations˜1, 3 and 2. Later, the satisfiability encoding that we will discuss next became the basis of [16] and [19].
For each non-edge $\{u,v\}\notin E$, we define Boolean variables $e_{uv}$, where $e_{uv}=\operatorname{True}$ means that $u$ and $v$ are merged, and thus have the same color. Likewise, $e_{uv}=\operatorname{False}$ means that the edge $\{u,v\}$ is added to the graph. For ease of notation, we extend the variables $e_{uv}$ also to proper edges $\{u,v\}\in E$. They are forced to $e_{uv}=\operatorname{False}$.
To ensure that each (partial) assignment of truth values corresponds to a valid graph in the Zykov tree, transitivity constraints (5) are required. They assert that if $u,v$ and $v,w$ have the same color, vertices $u$ and $w$ must also have the same color. With just the clauses from (5), any satisfying assignment corresponds to a leaf node in the Zykov tree but the number of colors is not yet restricted. To this end, Glorian et al. [16] use auxiliary variables $c_{v}$ for each vertex $v\in V$. In the clauses (6), $c_{v}=\operatorname{True}$ means that vertex $v$ needs a new color different from any previous vertex in the order given by $V=\{1,\dots,n\}$. Now, the sum over the $c_{v}$ variables counts the total number of used colors. The at-most-$k$-constraints (7) can be modeled as SAT in multiple ways, e.g., using cardinality networks [1] or the totalizer encoding [2]. This yields the following full (Zykov) encoding for the $k$-coloring problem.
@@ 2.3 IPASIR-UP Interface
We briefly repeat the main steps in a Conflict-Driven Clause Learning (CDCL) solver:
@@
The external propagator is notified by the SAT solver of newly assigned literals, new decision levels, and backtracking. It then has the possibility to modify the solving process using callbacks, as visualized in Figure˜2. For our application, we will use the following functionalities of the interface:
@@ 3 SAT-Based Propagator for the Zykov Encoding
In this section we adopt concepts of the propagator presented by Hébrard and Katsirelos [19] to the satisfiability setting using the interface discussed in Section˜2.3. This includes details on the transitivity constraints, using lower bounds for search tree pruning, and additional inferred propagations (Sections˜3.1, 3.2 and 3.3).
We further extend the SAT implementation with novel ideas and improvements, namely a better clique computation, modifications to the decision strategy, the design of an incremental bottom-up optimization that can reuse clauses and learned information, and the computation of fractional chromatic numbers during search (Sections˜3.4, 3.5, 3.6 and 3.7).
@@ 3.1 Transitivity Constraints
We follow the approach in [19] to avoid adding the cubic number of transitivity clauses (5) of the full encoding. Instead, the propagator ensures that $e_{uw}$ is set to $\operatorname{True}$ whenever $e_{uv}$ and $e_{vw}$ are set to $\operatorname{True}$. This is done efficiently by maintaining a graph $H$ that corresponds to the graph $G^{\prime}$, where $G^{\prime}$ is the graph in the Zykov tree obtained by following the merge operations and edge additions given by the currently assigned variables $e_{uv}$.
@@
These propagations are handled via cb_propagate, and the explanation clauses (5) are provided lazily using cb_add_reason, only should they be needed during conflict analysis. This is an advantage over the CP/SAT hybrid in [19], where transitivity clauses are immediately added to the encoding when the propagation is given. An important observation is that the propagator considers each variable only once per search path. This way, the transitivity propagation from the root node down to a full assignment only has complexity $O(n^{2})$[19]. In contrast, the constraint propagation from the root to a full assignment based on the full encoding of all transitivity constraints (5) has cubic complexity.
@@ 3.2 Lower Bound for Pruning
This lower bound-based pruning eliminates all solutions with more than $k$ colors. Thus, we do not need to add the clauses Equations˜6 and 7 explicitly an can avoid the auxiliary $c_{v}$ variables. If we find a lower bound that is larger than $k$, the assignment is pruned and the search is backtracked.
@@
To prune a partial assignment, we need to add a clause to the encoding that is unsatisfied under the current assignment. Since both the clique and Mycielskian bounds are based on a subgraph $H$, the explaining clause can be given by a description of the responsible subgraph:
@@
It forces that at least one edge of $E(H)$ must be true, i.e. $H$ is destroyed. The cb_add_external callback is used to add the clause to the encoding. Because it is unsatisfied at the node where the subgraph is present and the lower bound is found, this causes the desired backtracking. The clique heuristic is called at every node of the search tree, while the Mycielskian lower bound is only computed after backtracking and if, the gap between clique and upper bound is at most one. We did not observe a significant gain in pruning when the Mycielskian lower bound is computed regardless of the gap, in particular for bottom-up optimization. We thus choose to call the heuristic less frequently.
@@ 3.3 Inferred Propagations
In this section, we discuss another strategy. It does not prune the current partial assignment but infers necessary variable assignments. Given a subgraph with chromatic number $k$ and vertices $u$ and $v$, the propagator can determine that $e_{uv}$ must be assigned either $\operatorname{True}$ or $\operatorname{False}$, if otherwise a subgraph with chromatic number $k+1$ is formed.
@@
[…]
@@
Given a subgraph $H=(V^{\prime},E^{\prime})$ in $G=(V,E)$ with chromatic number $k$, if there exist two non-adjacent vertices $u$ and $v$ such that $V^{\prime}\subseteq(N(u)\cup N(v))$ then $G/(uv)$ has chromatic number at least $k+1$.
@@ 3.4 Improved Clique Computation
While very fast, the clique heuristic used in [19] (see Section˜3.2) is a greedy algorithm. It often does not produce near-optimum cliques. To improve the pruning potential of the propagator, we prefer an algorithm that produces larger cliques. Since the clique algorithm is called frequently in the propagator, keeping its running time small is essential.
@@ 3.5 Improved Decision Strategy
We use the reduction rule of dominated vertices to determine ?good? decisions and communicate them to the SAT solver. A vertex $u$ is dominated by a vertex $v$ if $N(u)\subseteq N(v)$ and we know that vertex $u$ can always take the same color as the more constrained vertex $v$. We can therefore merge $u$ into $v$ as there exists an optimal solution where the two vertices have the same color. Note that we can potentially also find an optimal solution where $u$ and $v$ have different colors. Thus, domination is weaker than the inferred propagations from the previous section. Therefore, we provide $e_{uv}=\operatorname{True}$ as a decision for the SAT solver using the cb_decide callback. The idea is that following this decision will lead to a part in the search tree that is more likely to contain a solution and thus should be checked first.